Renea is 5 years older than Paula
Renea is 7 years older than Peter
Between all 3 there total age is 21

Answer by checkley77(12844)

R=PA+5 OR PA=R-5
R=PE+7 OR PE=R-7
R+PA+PE=21
R+(R-5)+(R-7)=21
3R-12=21
3R=21+12
3R=33
R=33/3
R=11 YEARS OLD.
PA=11-5=6 YEARS OLD.
PE=11-7=4 YEARS OLD.
PROOF:
11+6+4=21
21=21